OpenStudy (anonymous):

Evaluate the expression by sketching a triangle. Sin(2cos^-1 3/5).

OpenStudy (anonymous):

let's start with \(\large cos^{-1}(\frac{3}{5})=y \). this means cosy = 3/5, correct? can you draw a triangle depicting this situation?

OpenStudy (anonymous):

going to take more than just a triangle i think

OpenStudy (anonymous):

we'll get to the double angle formula when we get there...

OpenStudy (anonymous):

you need to know that \(\sin(2\theta)=2\sin(\theta )\cos(\theta)\)

OpenStudy (anonymous):

cos=adj/hyp so the adj is 3 and the hypothenuse is 5

OpenStudy (anonymous):

using pythagorean theorem the opposite side will be 4

OpenStudy (anonymous):

but Idk what to do about the 2...

OpenStudy (anonymous):

How can I solve this?

OpenStudy (anonymous):

start with \[\sin(2\theta)=2\sin(\theta )\cos(\theta)\] and so \[\sin(\cos^{-1}(\frac{3}{5})=2\sin(\cos^{-1}(\frac{3}{5}))\times \cos(\cos^{-1}(\frac{3}{5}))\]

OpenStudy (anonymous):

typo meant \[\sin(2\cos^{-1}(\frac{3}{5})=2\sin(\cos^{-1}(\frac{3}{5}))\times \cos(\cos^{-1}(\frac{3}{5}))\]

OpenStudy (anonymous):

one number is obvious you know \(\cos(\cos^{-1}(\frac{3}{5}))=\frac{3}{5}\) that could be clear right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what is left is to find \(\sin(\cos^{-1}(\frac{3}{5}))\) which is identical to saying "if the cosine of the angle is \(\frac{3}{5}\) then what is the sine of the angle?

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of a triangle with the cosine of the angle \(\frac{3}{5}\) what you need is the opposite side, which by pythagoras or the famous 3-4-5 right triangle must be 4

OpenStudy (anonymous):



OpenStudy (anonymous):

now we see that the sine of that angle is \(\frac{4}{5}\) so we have all the numbers we need

OpenStudy (anonymous):

it is \[\sin(2\cos^{-1}(\frac{3}{5}))=2\sin(\cos^{-1}(\frac{3}{5}))\times \cos(\cos^{-1}(\frac{3}{5}))\] \[=2\times \frac{4}{5}\times \frac{3}{5}\] for your answer

OpenStudy (anonymous):

hope steps are clear

OpenStudy (anonymous):

yes, somewhat, thank you! =)